# Explicit Runge-Kutta Methods with MQ and IMQ-Radial Basis Functions

### v1

@@ 3.1 Two-Stage Third-Order Methods
where $\epsilon_{n}^{2}$ shape parameter and $\epsilon_{n}^{2}=0$ reduces the method to the classical form which is a classical Runge-Kutta method. Another form of the above method is
@@
By Taylor’s series expansion, the local truncation error becomes
$\tau$ $=f+\frac{h}{2}(f_{t}+ff_{u})+\frac{h^{2}}{3!}(f_{tt}+f_{tu}f+f(f_{tu}+ff_{uu})+f_{u}(f_{t}+ff_{u}))-b_{1}f$
$-b_{2}\left(f+c_{2}hf_{t}+\left(hfa_{21}-\frac{\epsilon_{n}^{2}a_{21}^{2}h^{2}u_{n}}{2}-\frac{\epsilon_{n}^{2}a_{21}^{2}h^{2}b_{2}c_{2}h^{3}}{2}\right)f_{u}\right)+\mathcal{O}(h^{3})$
$=\left(1-b_{1}-b_{2}\right)f+\left[\left(\frac{1}{2}-b_{2}c_{2}\right)f_{t}+\left(\frac{1}{2}-a_{21}b_{2}\right)f_{u}f\right]\,h$
$+\bigg{[}\left(\frac{1}{6}-\frac{1}{2}b_{2}c_{2}^{2}\right)f_{tt}+\left(\frac{1}{3}-a_{21}b_{2}c_{2}\right)f_{tu}f$
$+\left(\frac{1}{6}-\frac{1}{2}a_{21}^{2}b_{2}\right)f^{2}f_{uu}\bigg{]}h^{2}+\frac{(f_{t}+f_{u}f)f_{u}}{6}-\frac{\epsilon_{n}^{2}a_{21}^{2}b_{2}u_{n}f_{u}}{2}h^{2}+\mathcal{O}(h^{3}).$
The method will be third-order accurate if $a_{21}=c_{2}=\frac{2}{3}$, $b_{1}=\frac{1}{4}$, $b_{2}=\frac{3}{4}$ and
$\epsilon_{n}^{2}=\frac{f_{t}+ff_{u}}{u_{n}}.$
(b) IMQ
The two-stage IMQ-RBF Runge-Kutta method is of the form
$v_{n+1}=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}),$ (3.2)
with,
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f\left(t_{n}+c_{2}h,\sqrt{(1+\epsilon_{n}^{2}a_{21}^{2}h^{2})}hk_{1}+\frac{v_{n}}{\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}\right),$
where $\epsilon_{n}^{2}$ is shape parameter and equating to zero leads the above method to the classical RK-2 method. This modified method can be expressed by the augmented Butcher tableau,
$\begin{array}[]{c|cc|c}0&&\\ c_{2}&a_{21}&&\epsilon_{n}^{2}\\ \hline\cr&b_{1}&b_{2}\end{array}$ (3.3)
where $\epsilon_{n}^{2}$, $a_{ij}$, $b_{i}$, and $c_{i}$ carry conventional definition. This two-stage method can be rewritten as
$v^{(1)}$ $=\sqrt{(1+\epsilon_{n}^{2}a_{21}^{2}h^{2})}hk_{1}+\frac{v_{n}}{\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}},$
$v_{n+1}$ $=v_{n}\left(1-\frac{b_{1}}{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}\right)+\frac{b_{1}v^{(1)}}{\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}+b_{2}hf\left(t_{n}+c_{2}h,v^{(1)}\right).$
By Taylor’s series expansion, we have local truncation error as
$\tau$ $=\frac{u_{n+1}-u_{n}}{h}-(b_{1}k_{1}+b_{2}k_{2})$
$=f+\frac{h}{2}(f_{t}+ff_{u})+\frac{h^{2}}{3!}(f_{tt}+f_{tu}f+f(f_{tu}+ff_{uu})+f_{u}(f_{t}+ff_{u}))-b_{1}f$
$-b_{2}(f+c_{2}hf_{t}+\left(hfa_{21}-\frac{\epsilon_{n}^{2}a_{21}^{2}h^{2}u_{n}}{2}\right)+\mathcal{O}(h^{3})$
$=(1-b_{1}-b_{2})f+\left[\left(\frac{1}{2}-b_{2}c_{2}\right)f_{t}+\left(\frac{1}{2}-a_{21}b_{2}\right)f_{u}f\right]\,h$
$+\biggl{[}\left(\frac{1}{6}-\frac{1}{2}b_{2}c_{2}^{2}\right)f_{tt}+\left(\frac{1}{3}-a_{21}b_{2}c_{2}\right)f_{tu}f$
$+\left(\frac{1}{6}-\frac{1}{2}a_{21}^{2}b_{2}\right)f^{2}f_{uu}\biggl{]}h^{2}+(f_{t}+f_{u}f)f_{u}+\frac{\epsilon_{n}^{2}a_{21}^{2}b_{2}u_{n}f_{u}}{2}h^{2}+\mathcal{O}(h^{3}).$
The method will be third-order accurate if $a_{21}=c_{2}=\frac{2}{3}$, $b_{1}=\frac{1}{4}$, $b_{2}=\frac{3}{4}$ and
$\epsilon_{n}^{2}=\frac{f_{t}+ff_{u}}{u_{n}}.$
3.2 Three-stage forth-order methods
(a) MQ
$v_{n+1}=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3}),$ (3.4)
where,
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f\left(t_{n}+c_{2}h,(v_{n}+hk_{1}a_{21})\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right),$
$k_{3}$ $=f\left(t_{n}+c_{3}h,(v_{n}+h(k_{1}a_{31}+k_{2}a_{32}))\sqrt{1+\epsilon_{n3}^{2}a_{21}^{2}h^{2}}\right).$
The method can be written as
$v^{(1)}$ $=\left(v_{n}+hfa_{21}\right)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}},$
$v^{(2)}$ $=v_{n}\left(1-\frac{a_{31}}{a_{21}}\right)+\left[\frac{a_{31}v^{(1)}}{a_{21}\sqrt{1+\epsilon_{n}^{2}c_{2}^{2}h^{2}}}+ha_{32}f(t_{n}+c_{2}h,v^{(1)})\right]\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}},$
$v_{n+1}$ $=v_{n}\left[1-b_{1}h-\frac{b_{2}}{a_{32}h\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}+\frac{a_{31}}{a_{32}}\right]+v^{(1)}\left[\frac{b_{1}}{a_{21}\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}-\frac{b_{2}a_{31}}{a_{32}\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}\right]$
$\quad+\frac{v^{(2)}}{a_{32}h\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}+b_{3}f(t_{n}+c_{3}h,v^{(2)}).$
Using Taylor’s series expansion, the local truncation error derived as
$\tau_{n}$ $=\frac{u_{n+1}-u_{n}}{h}-(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3})$
$=(1-b_{1}-b_{2}-b_{3})+h\left(-a_{21}b_{2}ff_{u}-a_{31}b_{3}ff_{u}-a_{32}b_{3}ff_{u}-b_{2}c_{2}f_{t}-b_{3}c_{3}f_{t}+\frac{ff_{u}}{2}+\frac{f_{t}}{2}\right)$
$\quad+h^{2}\Bigg{(}-\frac{a_{21}^{2}b_{2}f^{2}f_{uu}}{2}-a_{21}a_{32}b_{3}ff_{u}^{2}-a_{21}b_{2}c_{2}ff_{ut}-\frac{a_{31}^{2}b_{3}f^{2}f_{uu}}{2}-\frac{a_{31}a_{32}b_{3}f^{2}f_{uu}}{2}$
$\quad-a_{31}b_{3}c_{3}ff_{ut}-\frac{a_{32}^{2}b_{3}f^{2}f_{uu}}{2}-a_{32}b_{3}c_{2}f_{t}f_{u}-a_{32}b_{3}c_{3}ff_{ut}-\frac{b_{2}c_{2}^{2}}{2}v_{\epsilon 2}f_{u}-\frac{b_{2}c_{2}^{2}f_{tt}}{2}-\frac{b_{3}c_{3}^{2}v_{\epsilon 3}f_{u}}{2}$
$\quad-\frac{b_{3}c_{3}^{2}f_{tt}}{2}+\frac{f^{2}f_{uu}}{6}+\frac{ff_{u}^{2}}{6}+\frac{ff_{ut}}{3}+\frac{f_{t}f_{u}}{6}+\frac{f_{tt}}{6}\Bigg{)}$
$\quad+h^{3}\Bigg{(}-\frac{a_{21}^{2}b_{2}f^{2}f_{uu}}{2}-a_{21}a_{32}b_{3}ff_{u}^{2}-a_{21}b_{2}c_{2}ff_{ut}-\frac{a_{31}^{2}b_{3}f^{2}f_{uu}}{2}-\frac{a_{31}a_{32}b_{3}f^{2}f_{uu}}{2}$
$\quad-a_{31}b_{3}c_{3}ff_{ut}-\frac{a_{32}^{2}b_{3}f^{2}f_{uu}}{2}-a_{32}b_{3}c_{2}f_{t}f_{u}-a_{32}b_{3}c_{3}ff_{ut}-\frac{b_{2}c_{2}^{2}}{2}v_{\epsilon 2}f_{u}-\frac{b_{2}c_{2}^{2}f_{tt}}{2}-\frac{b_{3}c_{3}^{2}v_{\epsilon 3}f_{u}}{2}$
$\quad-\frac{b_{3}c_{3}^{2}f_{tt}}{2}+\frac{f^{2}f_{uu}}{6}+\frac{ff_{u}^{2}}{6}+\frac{ff_{ut}}{3}+\frac{f_{t}f_{u}}{6}+\frac{f_{tt}}{6}\Bigg{)}+\mathcal{O}(h^{3})$
@@ 3.2 Three-stage forth-order methods
are imposed the above local truncation error term vanishes. There are three cases where different parameter values makes the method forth order accurate.
@@
    $a_{21}=c_{2}=\frac{1}{3},a_{31}=-\frac{5}{12},a_{32}=\frac{5}{4},c_{3}=\frac{5}{6},b_{1}=\frac{1}{10},b_{2}=\frac{1}{2},b_{3}=\frac{2}{5}.$
    $\epsilon^{2}_{n2}$ $=\frac{ff_{t}f_{uu}-3ff_{u}^{3}-ff_{u}f_{tu}-3f_{t}f_{u}^{2}+f_{t}f_{tu}-f_{tt}f_{u}}{vff_{uu}-2vf_{u}^{2}+vf_{tu}+ff_{u}},$
    $\epsilon^{2}_{n3}$ $=-\frac{1}{5}\epsilon^{2}_{n2}.$
@@
    $a_{21}=c_{2}=1,a_{31}=\frac{1}{4},a_{32}=\frac{1}{4},c_{3}=\frac{1}{2},b_{1}=\frac{1}{6},b_{2}=\frac{1}{6},b_{3}=\frac{2}{3}.$
    $\epsilon^{2}_{n2}$ $=\frac{ff_{t}f_{uu}+ff_{u}^{3}-ff_{u}f_{tu}+f_{t}f_{u}^{2}+f_{t}f_{tu}-f_{tt}f_{u}}{vff_{uu}+2vf_{u}^{2}+vf_{tu}+ff_{u}},$
    $\epsilon^{2}_{n3}$ $=-\epsilon^{2}_{n2}.$
@@
    $a_{21}=c_{2}=\frac{1}{2},a_{31}=0,a_{32}=c_{3}=\frac{3}{4},c_{3}=1,b_{1}=\frac{2}{9},b_{2}=\frac{1}{3},b_{3}=\frac{4}{9}.$
    $\epsilon^{2}_{n2}$ $=\frac{-\frac{1}{3}f^{3}f_{uuu}-f^{2}f_{tuu}-4ff_{u}^{3}-ff_{ttu}-4f_{t}f_{u}^{2}-\frac{1}{3}f_{ttt}}{vff_{uu}-4vf_{u}^{2}+vf_{tu}+ff_{u}}$
    $\epsilon^{2}_{n3}$ $=-\frac{1}{3}\epsilon^{2}_{n2}.$
(b) IMQ
$v_{n+1}=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3}),$ (3.5)
where,
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f\left(t_{n}+c_{2}h,\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}hf_{n}+\frac{v_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\right),$
$k_{3}$ $=f\left(t_{n}+c_{3}h,\sqrt{(1+\epsilon_{n3}^{2}c_{3}^{2}h^{2})}h(a_{31}k_{1}+a_{32}k_{2})+\frac{v_{n}}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\right).$

### v2

@@ 3.1 Two-Stage Third-Order Methods
where $\epsilon_{n}^{2}$ is a shape parameter and $\epsilon_{n}^{2}=0$ reduces the method to the classical form of Runge-Kutta method. Alternative form of the (3.3) is the following.
$\begin{cases}\begin{aligned} v^{(1)}&=\sqrt{(1+\epsilon_{n}^{2}a_{21}^{2}h^{2})}\,\,(v_{n}+hk_{1}a_{21}),\\ v_{n+1}&=v_{n}\left(1-\dfrac{b_{1}}{a_{21}h}\right)+v^{(1)}\dfrac{b_{1}}{\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}+f(t_{n}+c_{2}h,v^{(1)})b_{2}h.\end{aligned}\end{cases}$
Using a Taylor series expansion, the local truncation error is expressed in powers of $h.$ Specifically, we expand $k_{2}$ using the two-variable Taylor series about the point $(t_{n},u_{n}),$ while the derivative of $u_{n}$ is obtained from the relation (1.1) satisfied by $u_{n}.$ Assuming $v_{n}=u_{,}n$ the resulting expression represents the error produced after a single time step for (3.3) is
$\tau$ $=\frac{u_{n+1}-u_{n}}{h}-(b_{1}k_{1}+b_{2}k_{2}),$
$=\dfrac{1}{h}\left(u_{n}+hu_{n}^{\prime}+\frac{h^{2}}{2}u_{n}^{\prime\prime}+\frac{h^{3}}{3!}u_{n}^{\prime\prime\prime}+\cdots-u_{n}\right)$
$\qquad-\left(b_{1}f(t_{n},v_{n})+b_{2}f\left(t_{n}+c_{2}h,(v_{n}+hk_{1}a_{21})\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}\right)\right),$
$=\left(f+\frac{h}{2}(f_{t}+ff_{u})+\frac{h^{2}}{3!}(f_{tt}+f_{tu}f+f(f_{tu}+ff_{uu})+f_{u}(f_{t}+ff_{u}))\right)(t_{n},u_{n})$
$\qquad-b_{1}f(t_{n},u_{n})-b_{2}\left(f+c_{2}hf_{t}+\left(hfa_{21}-\frac{\epsilon_{n}^{2}a_{21}^{2}h^{2}u_{n}}{2}-\frac{\epsilon_{n}^{2}a_{21}^{2}h^{2}b_{2}c_{2}h^{3}}{2}\right)f_{u}\right)(t_{n},u_{n})$
$\qquad+\mathcal{O}(h^{3}),$
$=\left(1-b_{1}-b_{2}\right)f(t_{n},u_{n})+h\left[\left(\frac{1}{2}-b_{2}c_{2}\right)f_{t}+\left(\frac{1}{2}-a_{21}b_{2}\right)f_{u}f\right](t_{n},u_{n})$
$\qquad+h^{2}\bigg[\left(\frac{1}{6}-\frac{1}{2}b_{2}c_{2}^{2}\right)f_{tt}+\left(\frac{1}{3}-a_{21}b_{2}c_{2}\right)f_{tu}f+\left(\frac{1}{6}-\frac{1}{2}a_{21}^{2}b_{2}\right)f^{2}f_{uu}\bigg](t_{n},u_{n})$
$\qquad+h^{2}\left(\frac{(f_{t}+f_{u}f)f_{u}}{6}-\frac{\epsilon_{n}^{2}a_{21}^{2}b_{2}u_{n}f_{u}}{2}\right)(t_{n},u_{n})+\mathcal{O}(h^{3}).$
The method will maintain second-order accuracy under the consistency and order conditions (2.5), (2.2), and (2.3). In addition, the method will be third order accurate if $a_{21}=c_{2}=\frac{2}{3}$, $b_{1}=\frac{1}{4}$, $b_{2}=\frac{3}{4}$ and
$\epsilon_{n}^{2}=\dfrac{f_{t}+ff_{u}}{u_{n}},$
as the truncation error terms vanish upto $\mathcal{O}(h^{3}).$
(b) IMQ-RBF RK scheme: The two-stage IMQ-RBF Runge-Kutta method is of the form
$\begin{cases}\begin{aligned} v_{n+1}&=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}),\\ k_{1}&=f(t_{n},v_{n}),\\ k_{2}&=f\left(t_{n}+c_{2}h,\sqrt{(1+\epsilon_{n}^{2}a_{21}^{2}h^{2})}hk_{1}+\frac{v_{n}}{\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}\right),\end{aligned}\end{cases}$ (3.4)
where $\epsilon_{n}^{2}$ is the shape parameter and making it to zero, then method (3.4) leads to the classical RK-2 method. This two-stage method can be rewritten as
$\begin{cases}\begin{aligned} v^{(1)}&=v_{n}\dfrac{1}{\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}+\sqrt{(1+\epsilon_{n}^{2}a_{21}^{2}h^{2})}hk_{1},\\ v_{n+1}&=v_{n}\left(1-\frac{b_{1}}{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}\right)+v^{(1)}\dfrac{b_{1}}{\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}+b_{2}hf\left(t_{n}+c_{2}h,v^{(1)}\right).\end{aligned}\end{cases}$
Using Taylor’s series expansion, we have local truncation error as
$\begin{aligned} \tau&=(1-b_{1}-b_{2})f(t_{n},u_{n})+h\left[\left(\frac{1}{2}-b_{2}c_{2}\right)f_{t}+\left(\frac{1}{2}-a_{21}b_{2}\right)f_{u}f\right](t_{n},u_{n})\\ &\qquad+h^{2}\biggl[\left(\frac{1}{6}-\frac{1}{2}b_{2}c_{2}^{2}\right)f_{tt}+\left(\frac{1}{3}-a_{21}b_{2}c_{2}\right)f_{tu}f\\ &\qquad+\left(\frac{1}{6}-\frac{1}{2}a_{21}^{2}b_{2}\right)f^{2}f_{uu}\biggl](t_{n},u_{n})+h^{2}\bigg[(f_{t}+f_{u}f)f_{u}+\frac{\epsilon_{n}^{2}a_{21}^{2}b_{2}u_{n}f_{u}}{2}\bigg](t_{n},u_{n})+\mathcal{O}(h^{3}).\end{aligned}$
The method will maintain second-order accuracy under the consistency and order conditions (2.5), (2.2), and (2.3). In addition, the method will be third order accurate if The method will be third-order accurate if $a_{21}=c_{2}=\frac{2}{3}$, $b_{1}=\dfrac{1}{4}$, $b_{2}=\frac{3}{4}$ and $\epsilon_{n}^{2}=-\dfrac{f_{t}+ff_{u}}{u_{n}}.$
In summary, these two-stage third-order methods can be expressed by the augmented Butcher tableau as follows;
$\begin{array}[]{c|cc|c}0&&\\ c_{2}=\dfrac{2}{3}&a_{21}=\dfrac{2}{3}&&\textit{MQ:}\,\epsilon_{n}^{2}=\dfrac{f_{t}+ff_{u}}{u_{n}}\\ \hline\cr&b_{1}=\dfrac{1}{4}&b_{2}=\dfrac{3}{4}\end{array}$ (3.5)
$\begin{array}[]{c|cc|c}0&&\\ c_{2}=\dfrac{2}{3}&a_{21}=\dfrac{2}{3}&&\textit{IMQ:}\,\epsilon_{n}^{2}=-\dfrac{f_{t}+ff_{u}}{u_{n}}\\ \hline\cr&b_{1}=\dfrac{1}{4}&b_{2}=\dfrac{3}{4}\end{array}$ (3.6)
3.2 Three-stage fourth-order methods
In this subsection, we develop three-stage fourth-order MQ and IMQ-RBF based Runge-Kutta schemes and perform the turncation error analysis.
(a) MQ-RBF RK scheme: The three-stage MQ-RBF Runge-Kutta scheme is of the form
$\begin{cases}\begin{aligned} v_{n+1}&=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3}),\\ k_{1}&=f(t_{n},v_{n}),\\ k_{2}&=f\left(t_{n}+c_{2}h,(v_{n}+hk_{1}a_{21})\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right),\\ k_{3}&=f\left(t_{n}+c_{3}h,(v_{n}+h(k_{1}a_{31}+k_{2}a_{32}))\sqrt{1+\epsilon_{n3}^{2}a_{21}^{2}h^{2}}\right).\end{aligned}\end{cases}$ (3.7)
The method can be written in alternative form as
$\begin{cases}\begin{aligned} v^{(1)}&=\left(v_{n}+hfa_{21}\right)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}},\\ v^{(2)}&=v_{n}\left(1-\dfrac{a_{31}}{a_{21}}\right)+\left[\dfrac{a_{31}}{a_{21}\sqrt{1+\epsilon_{n}^{2}c_{2}^{2}h^{2}}}v^{(1)}+ha_{32}f(t_{n}+c_{2}h,v^{(1)})\right]\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}},\\ v_{n+1}&=v_{n}\left[1-b_{1}h-\frac{b_{2}}{a_{32}h\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}+\frac{a_{31}}{a_{32}}\right]\\ &\qquad+v^{(1)}\left[\dfrac{b_{1}}{a_{21}\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}-\frac{b_{2}a_{31}}{a_{32}\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}}\right]+\dfrac{v^{(2)}}{a_{32}h\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\\ &\qquad\qquad+b_{3}f(t_{n}+c_{3}h,v^{(2)}).\end{aligned}\end{cases}$
Upon Taylor’s series expansion for (3.7), the local truncation error can be derived as
$\tau_{n}$ $=\frac{u_{n+1}-u_{n}}{h}-(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3}),$
$=(1-b_{1}-b_{2}-b_{3})f+h\left[ff_{u}\left(\frac{1}{2}-a_{21}b_{2}-a_{31}b_{3}-a_{32}b_{3}\right)+f_{t}\left(\frac{1}{2}-b_{3}c_{3}-b_{2}c_{2}\right)\right]$
@@ 3.2 Three-stage fourth-order methods
$\quad\left.\left.-a_{31}b_{3}c_{3}-a_{32}b_{3}c_{3}\right)+f_{t}f_{u}\left(\frac{1}{6}-a_{32}b_{3}c_{2}\right)+f_{tt}\left(\frac{1}{6}-\frac{b_{2}c_{2}^{2}}{2}-\frac{b_{3}c_{3}^{2}}{2}\right)-\left(b_{2}c_{2}^{2}\epsilon_{n2}^{2}+b_{3}c_{3}^{2}\epsilon_{n3}^{2}\right)u_{n}f_{u}\right]$
[…]
$\quad\left.-a_{31}a_{32}b_{3}c_{2}-a_{32}^{2}b_{3}c_{2}\right)+ff_{utt}\left(\frac{1}{8}-\frac{a_{31}b_{3}c_{3}^{2}}{2}-\frac{a_{32}b_{3}c_{3}^{2}}{2}\right)+f_{tt}f_{u}\left(\frac{1}{24}-\frac{a_{32}b_{3}c_{2}^{2}}{2}\right)-f_{t}f_{ut}\left(\frac{1}{8}\right$.
$\quad\left.\left.-a_{32}b_{3}c_{2}c_{3}\right)+f_{ttt}\left(\frac{1}{24}-\frac{b_{2}c_{2}^{3}}{6}-\frac{b_{3}c_{3}^{3}}{6}\right)-\frac{a_{32}b_{3}c_{2}^{2}u_{n}f_{u}^{2}\epsilon_{2n}^{2}}{2}-\frac{a_{31}b_{3}c_{3}^{2}\epsilon_{3n}^{2}ff_{u}}{2}-\frac{a_{31}b_{3}c_{3}^{2}\epsilon_{3n}^{2}u_{n}ff_{u}}{2}\right$.
$\quad\left.-\frac{a_{32}b_{3}c_{3}^{2}\epsilon_{3n}^{2}u_{n}ff_{uu}}{2}-\frac{a_{32}b_{3}c_{3}^{2}ff_{u}\epsilon_{3n}^{2}}{2}-\frac{b_{2}c_{2}^{3}u_{n}\epsilon_{2n}^{2}f_{ut}}{2}-\frac{b_{3}c_{3}^{3}u_{n}\epsilon_{3n}^{2}f_{ut}}{2}\right]+\mathcal{O}(h^{4}).$
@@
are imposed the above local truncation error term vanishes. There are three cases where different parameter values makes the method forth-order accurate.
@@
    $\begin{aligned} &a_{21}=c_{2}=\frac{1}{3},a_{31}=-\frac{5}{12},a_{32}=\frac{5}{4},c_{3}=\frac{5}{6},b_{1}=\frac{1}{10},b_{2}=\frac{1}{2},b_{3}=\frac{2}{5}.\\ &\epsilon^{2}_{n2}=\frac{ff_{t}f_{uu}-3ff_{u}^{3}-ff_{u}f_{tu}-3f_{t}f_{u}^{2}+f_{t}f_{tu}-f_{tt}f_{u}}{zvff_{uu}-2vf_{u}^{2}+vf_{tu}+ff_{u}}\\ &\epsilon^{2}_{n3}=-\frac{1}{5}\epsilon^{2}_{n2}.\end{aligned}$
@@
    $\begin{aligned} &a_{21}=c_{2}=1,a_{31}=\frac{1}{4},a_{32}=\frac{1}{4},c_{3}=\frac{1}{2},b_{1}=\frac{1}{6},b_{2}=\frac{1}{6},b_{3}=\frac{2}{3}.\\ &\epsilon^{2}_{n2}=\frac{ff_{t}f_{uu}+ff_{u}^{3}-ff_{u}f_{tu}+f_{t}f_{u}^{2}+f_{t}f_{tu}-f_{tt}f_{u}}{vff_{uu}+2vf_{u}^{2}+vf_{tu}+ff_{u}}\\ &\epsilon^{2}_{n3}=-\epsilon^{2}_{n2}.\end{aligned}$
@@
    $\begin{aligned} &a_{21}=c_{2}=\frac{1}{2},a_{31}=0,a_{32}=c_{3}=\frac{3}{4},c_{3}=1,b_{1}=\frac{2}{9},b_{2}=\frac{1}{3},b_{3}=\frac{4}{9}.\\ &\epsilon^{2}_{n2}=\frac{-\frac{1}{3}f^{3}f_{uuu}-f^{2}f_{tuu}-4ff_{u}^{3}-ff_{ttu}-4f_{t}f_{u}^{2}-\frac{1}{3}f_{ttt}}{vff_{uu}-4vf_{u}^{2}+vf_{tu}+ff_{u}}\\ &\epsilon^{2}_{n3}=-\frac{1}{3}\epsilon^{2}_{n2}.\end{aligned}$
(b) IMQ-RBF RK scheme:
The three-stage IMQ-RBF Runge-Kutta scheme is of the form
$\begin{cases}\begin{aligned} v_{n+1}&=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3}),\\ k_{1}&=f(t_{n},v_{n}),\\ k_{2}&=f\left(t_{n}+c_{2}h,\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}ha_{21}f_{n}+\frac{v_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\right),\\ k_{3}&=f\left(t_{n}+c_{3}h,\sqrt{(1+\epsilon_{n3}^{2}c_{3}^{2}h^{2})}h(a_{31}k_{1}+a_{32}k_{2})+\frac{v_{n}}{\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}\right).\end{aligned}\end{cases}$ (3.8)